# Continual Learning of Large Language Models: A Comprehensive Survey

### v1

@@ 5 Evaluation Protocols and Datasets
In this section, we introduce the evaluation protocols and datasets for continually learning large language models. In Section 5.1, we discuss common continual learning evaluation metrics adapted for this context, along with metrics designed specifically for continual LLMs. Then, in Section 5.2, we outline the datasets available for each discussed topic.
5.1 Evaluation Protocols
In the realm of conventional continual learning, where task streams take the form of classification, many metrics rely on the concept of Accuracy Matrix [174, 260]. Extending this notion to the context of continually learning LLMs, we introduce the Performance Matrix ${\bm{P}}\in\mathbb{R}^{T\times T}$, where $T$ represents the total number of training stages. Each entry of ${\bm{P}}$ corresponds to a performance metric evaluated on the models, such as perplexity on pre-training data [127, 46, 89], zero-shot/few-shot evaluation metrics on downstream data without fine-tuning [52, 311, 10, 63, 215, 245], fine-tuned accuracies on downstream tasks [6, 231, 46, 120], and probing accuracies from fine-tuning add-on components evaluated on downstream tasks [274, 183, 365]. In ${\bm{P}}$, $P_{i,j}$ denotes the model’s performance after training on task $i$ and evaluating on task $j$. With this Performance Matrix definition, we introduce the primary evaluation protocols widely adopted.
Overall Performance (OP). The Overall Performance (OP) [133, 357, 351] is a natural extension of the concept of Average Accuracy [174, 260]. The OP measured up until training stage $t$ is the average performance of the model trained right after the stage $t$. Denote it as $\operatorname{OP}_{t}$ and we have:
$\operatorname{OP}_{t}$ $\triangleq\tfrac{1}{t}\sum_{i=1}^{t}P_{t,i}.$ (14)
As noted in [260], the OP corresponds to the primary optimization objective defined in Definition 2.5, 2.6, and 2.7. In much of the continual learning literature, once all $T$ tasks are completed, the final $\operatorname{OP}$ ($\operatorname{OP}_{T}$) is reported, with the subscript T often omitted for brevity. In some works, OP is weighted by the importance of tasks $\widetilde{\operatorname{OP}}\triangleq\tfrac{1}{T}\sum_{i=1}^{T}w_{i}P_{t,i}$, where $w_{i}=N_{i}/\sum_{j=1}^{T}N_{j}$ represents the ratio of data. In some literature, $\widetilde{\operatorname{OP}}$ is referred to as “example accuracy” [47], “whole accuracy” [268] or “edit success rate” in CMR [96].
$F_{t}$ $\triangleq\tfrac{1}{t-1}\sum_{j=1}^{t-1}\left[\max_{l\in[t-1]}\{P_{l,j}-P_{t,j}\}\right].$ (15)
Typically, researchers report the average forgetting $F=F_{T}$ at the end of the entire training process. Forgetting quantifies the impact of learning new tasks on previously acquired knowledge. Ideally, a robust continual learning framework should achieve Backward Transfer (BWT), where learning new tasks enhances performance on prior tasks. This enhancement is typically measured by negating the forgetting, thus indicating an improvement in performance on earlier tasks. The concepts of Forgetting and Backward Transfer underpin various evaluation metrics, such as knowledge retention [127], performance on unchanged knowledge [119], average increased perplexity (AP+[232], and test and edit retention rate in CMR [96].
$\operatorname{FWT}_{t}$ $\triangleq\tfrac{1}{t-1}\sum_{i=2}^{t}P_{i-1,i}-b_{i},$ (16)
where $b_{i}$ is the baseline performance of the model evaluated on task $i$ before undergoing continual learning. Strictly speaking, the definition of $b_{i}$ is not the same as defined in the previous work [174, 260], where it is used to denote the performance of a random initialization of the model. Additionally, we extend the notation of forward transfer in the vertical direction to represent the performance improvement on downstream tasks resulting from domain-adaptive pre-training (see Table 2). Forward Transfer is alternatively referred to as temporal generalization [127] or knowledge transfer [145] in some literature.
@@ 5.1 Evaluation Protocols
LAnguage Model Analysis (LAMA). LAnguage Model Analysis (LAMA) is an evaluation framework designed to probe the world knowledge embedded in language models [225]. It converts each world fact into a cloze statement, which is then inputted into the language models to predict the correct answer. LAMA has been extended for continual pre-training, particularly for those under the temporal shifts [119, 120]. In CKL, three LAMA benchmarks are constructed for different dimensions: InvariantLAMA assesses knowledge retention on time-invariant facts, UpdatedLAMA focuses on knowledge update, and NewLAMA evaluates knowledge acquisition [120].
@@
Forgotten / (Updated + Acquired) Ratio (FUAR). As the performance of a pre-trained LLM is decomposed into a fine-grained set in CKL [120], OP becomes a too general metric and cannot accurately reflect the balance and trade-offs of the model’s behavior. To address this issue, CKL proposes a joint evaluation metric FUAR (Forgotten / (Updated + Acquired) Ratio) for continual pre-training. A FUAR value of 1 represents an equal trade-off between the knowledge forgetting and knowledge learning: for each piece of updated or acquired knowledge, one piece of time-invariant knowledge is forgotten on average. A FUAR less than 1 suggests high learning efficacy, where more than one piece of knowledge is acquired at the expense of forgetting one piece of time-invariant knowledge.
@@
X-Delta. In TRACE [292], the authors propose a set of "X-Delta" metrics for continual instruction tuning, quantifying the forward transfer on specific abilities of LLMs. Let’s denote a set of $M$ datasets $\{X_{1},X_{2},\cdots,X_{M}\}$ for task X. The baseline performances of the pre-trained LLM evaluated on these tasks are denoted as $\{b_{1}^{X},\cdots,b_{M}^{X}\}$. The model undergoes continuous fine-tuning on a different set of tasks, distinct from those used for evaluation. Throughout the sequential training process, the performance of the model after learning task $t$ on evaluation tasks $X_{i}$ is $R_{t,i}^{X}$. The X-Delta $\Delta R_{t}^{X}$ after learning task $t$ is defined as:
$\Delta R_{t}^{X}$ $\triangleq\tfrac{1}{M}\sum_{m=1}^{M}(R_{t,i}^{X}-b_{i}^{X}).$ (17)
@@
In the public TRACE benchmark, the authors construct three sets of evaluation tasks to benchmark the ability of LLMs, including general ability, instruction following, and safety [292].
@@
NLG Score. In continual model alignment, three prominent metrics used to evaluate different aspects of Natural language generation (NLG) are BLEU-4 [220], METEOR [14], and ROUGE-L [163]. BLEU-4[220], designed for machine translation (MT), evaluates the precision of n-grams between the machine-generated and reference texts, focusing especially on four-word sequences to gauge fluency and adequacy. METEOR [14] also targets MT but aims to improve correlation with human judgment by considering synonyms and stemming, thus providing a more nuanced assessment of translation quality. On the other hand, ROUGE-L [163] is commonly applied in summarization tasks, assessing the longest common subsequence between the generated summary and a set of reference summaries, effectively measuring the recall of essential content. Each metric has its strengths and is tailored to specific kinds of language processing tasks, reflecting different dimensions of text generation quality.
rPMS. The reference PM score (rPMS) quantifies the alignment of a model’s outputs with human preferences as captured by a reference Preference Model (PM). Specifically, for a given task $t$, the rPMS is defined as:
$\text{rPMS}_{t}=\text{rPM}(M_{t},D_{\text{test},t}),$ (18)
where $M_{t}$ is the model being evaluated, $D_{\text{test},t}$represents the test dataset for task $t$, and rPMS is the function implemented by the reference PM that scores model outputs based on their alignment with human preferences. Higher rPMS values indicate that the model $M_{t}$ aligns closely with human preferences, suggesting effective learning and retention of the desired behaviors across tasks in a continual learning scenario. The metric is crucial for evaluating the extent to which models can maintain or improve performance relative to human preference standards without substantial forgetting over time.
5.2 Datasets
@@
In this section, we provide a comprehensive review of the datasets available for benchmarking continual LLMs, as illustrated in Table 4. We intentionally exclude datasets used for domain-adaptive pre-training LLMs in vertical domains such as legal, medical, and financial, unless they are specifically designed for continual domain-adaptive pre-training. Furthermore, we omit datasets used in general continual fine-tuning, as they have already been extensively studied in existing works [22, 132].
@@
Datasets for Continual Pre-Training (CPT) and Domain Adaptive Pre-Training (DAP). Current research lacks a widely recognized benchmark for evaluating continual pre-training LLMs under temporal shifts. TimeLMs utilizes a series of Twitter corpora collected until 2022, sequentially pre-training RoBERTa models quarterly [175]. CC-RecentNews, adopted as unlabeled pre-training data for LMs in CKL [120], consists of recent news and serves as a single-stage dataset. Additionally, CKL introduces InvariantLAMA, NewLAMA, and UpdatedLAMA to assess the principles of continual knowledge learning. TWiki, a dataset derived from the articles of Wikipedia between August and December 2021, is curated and cleaned in TemporalWiki [119]. This dataset facilitates the exploration of incremental learning by providing the Diffsets between neighboring snapshots. For works that study the content-level distributional shifts in CPT and DAP, researchers often resort to a similar set of publicly available datasets [172, 325, 211] to construct their own test beds for continual learning algorithms. The DAPT dataset, developed by [92], comprises four domains: BioMed and Computer Science from S2ORC [172], News from [346], and Reviews from [101]. In DAPT’s original study, each domain undergoes individual domain adaptive pre-training stages to demonstrate the universality of DAP’s effectiveness. Subsequent works, such as ELLE [232] and Recyclable Tuning [231], follow suit by employing these domains for multi-stage CPT. DEMix [91] presents another large-scale dataset, featuring eight semantic domains with over 73.8 billion tokens. Alongside the training set, it includes eight additional datasets for validating the generalization ability of LLMs. On a smaller scale, CPT [131] and DAS [134] datasets consist of four and eight domains, respectively, with approximately 3.12 million examples and a size of 4.16GB each. These datasets are constructed similarly to the aforementioned ones.
@@
Table 4: Summary of the existing benchmarks publicly available for Continual Learning LLMs. In the column of Name, we use the superscript “” to denote the lack of the dataset name and the name shown is that of the original paper. In this table, we deliberately omit the datasets used for domain-adaptive pre-training the vertical LLMs, as their main focus of development is not on continual learning. We also omit the datasets used for general continual fine-tuning, as they are extensively discussed in other existing surveys [22, 132].
@@
TimeLMs [175] CPT Temporal Social Media 8 #Examples: 123.86M Tweets [175] code
CC-RecentNews [120] CPT Temporal News 1 #Tokens: $\sim$168M Web [120] code
@@
TWiki [119] CPT Temporal General Knowledge 5 #Tokens: 4.7B Wikipedia [119] code
@@
DAPT [92] CPT DAP Content Multi-Domain 4 Size: 160GB BioMed [172], CS [172], News [346], Reviews [101] [92] [231] [232] code
@@
CPT [131] CPT Content Multi-Domain 4 #Examples: 3.12M Yelp [325], S2ORC [172], AG-News [356] [131] code
@@
DEMix [91] CPT Content Multi-Domain 8 #Tokens: 73.8B 1B [39], CS [172], Legal [34], Med [172] WebText [83], RealNews [346], Reddit [17], Reviews [211] [91] code
@@
DAS [134] CPT DAP Content Multi-Domain 6 Size: 4.16GB Yelp [325], Reviews [211], Papers [172], PubMed [134] code
@@
SuperNI [295] CIT Content Mutli-Domain 16 #Tasks: 1616 #Examples: $\sim$5M GitHub [358, 294] code
@@
CITB [358] CIT Content Mutli-Domain 19 #Tasks: 38 SuperNI [295] [358] code
@@
CoIN [41] CIT Content Multi-Domain 8 #Examples: $\sim$1.14M RefCOCO [130],RefCOCO+ [190],RefCOCOg [190] ImageNet [64], VQAv2 [85], ScienceQA [179] TextVQA  [261], GQA [118], VizWiz  [90], OCR-VQA [200] [41] code
@@
TRACE [292] CIT Content Mutli-Domain 8 #Examples: 56,000 ScienceQA [179], FOMC [257], MeetingBank [111] C-STANCE [359], 20Minuten [135], CodeXGLUE [180], NumGLUE[202] [292] code
@@
NATURAL-INSTRUCTION [201] CIT Content Mutli-Domain 6 #Examples: 193k CosmosQA [114], DROP [73], Essential-Terms [137] MCTACO [368], MultiRC [136], QASC [138] Quoref[59] , ROPES [164] , Winogrande [248] [201] code
@@
IMDB [188] CMA Content Social Media 1 Size: 217.35 MB IMDB [351] code
@@
HH-RLHF [13] CMA Content General Knowledge 1 Size: 28.1 MB Human Feedback [351] code
@@
Reddit TL;DR [284] CMA Content Social Media 2 Size: 19.6 GB Reddit [351, 352] code
@@
Common Sense QA [165] Reading Comprehension [165] Translation [165] CMA Content Multi-Domain 6 #Examples: $\sim$ 41.16M ARC Easy and Challenge [51], Race [144], PIQA [23] SQuAD [236], DROP [73] WMT 2014 French to English [24] [165] see sources
@@
FEVER [278] CMR Content General Knowledge 1 #Examples: 420k Wikipedia [60, 98] code
@@
VitaminC [253] CMR Content General Knowledge 1 #Examples: 450k Wikipedia [204] code
zsRE [147] CMR Content General Knowledge 1 #Examples: 120M Wikireading [106] [98, 196, 197, 97, 96, 58] code
@@
T-rex [76] CMR Content General Knowledge 1 #Examples: 11M Dbpedia abstracts [28] [150, 70] code
@@
NQ [142] CMR Content General Knowledge 1 #Examples: 320k Google queries, Wikipedia [96] code
@@
CounterFact [196] CMR Content General Knowledge 1 #Examples: 22k zsRE [147] [196, 339, 108, 58] code
@@
SCOTUS [36] CMR Temporal Law 1 #Examples: 9.2k Supreme Court Database [96] code
@@
Datasets for Continual Instruction Tuning. Measuring the effectiveness of CIT is crucial, particularly because traditional evaluation metrics may not be suitable for LLMs: many of them are overly simplistic and fail to comprehensively assess the model’s ability to learn continually. New benchmarks and metrics are required to evaluate both the retention of old knowledge and the integration of new instructions. TRACE [292] stands as a continual learning benchmark designed specifically for LLMs, encompassing diverse tasks such as multilingual capabilities, code generation, and mathematical reasoning. CITB [358] represents another benchmark for CIT, incorporating both learning and evaluation protocols. It in addition demonstrates that replay generally yields the best performance across all methods. CoIN [41] extends the benchmark to MLLMs, incorporating a balanced and diverse set of instructions from vision-language datasets.
@@
Datasets for Continual Model Refinement. Most datasets for continual model refinement can be categorized into two types [192]: fact checking and question answering. For fact checking, models are asked to verify the truthfulness of certain claims, typically modeled as a classification task. Key datasets include FEVER [278] (used by [60, 98]) and VitaminC [253] (used by [204]), both sourced from Wikipedia. For question answering, models are tasked with providing specific answers instead of choices. Zero-shot Relation Extraction (zsRE) [147] is the most widely employed dataset for this purpose  [98, 196, 197, 97, 96, 58], alongside Natural Questions (NQ)  [142] and T-rex  [76]. [196] adapted zsRE with additional counterfactuals to create the more challenging CounterFact dataset, used by  [339, 108, 58]. Beyond these two categories, SCOTUS [36] is also utilized [96] in the assessment of continual model refinement through a document classification task for U.S. Supreme Court cases into 11 topics.
@@
Datasets for Continual Model Alignment. In the domain of reinforcement learning with human feedback (RLHF), several datasets are commonly employed across different studies to evaluate the adaptation and effectiveness of models under varying scenarios and continuous learning conditions. The IMDB [188] and HH-RLHF [13] dataset, as introduced in [351] within their study on continual learning through optimal policy fitting, leverages data gathered from interactive RL scenarios to model human preferences dynamically. Similarly, the Reddit TL;DR dataset [284] used by [352, 351] is focused on text summarization, providing a robust platform for testing the longevity and adaptability of learning algorithms under evolving conditions. Lastly, Common Sense QA [51, 144, 23], Reading Comprehension [236, 73], and Translation [24], which are utilized in [165] are selected to assess the challenges of aligning RL agents with human expectations without incurring significant performance penalties. Each of these datasets is pivotal in advancing the understanding of continual learning and the interplay between human feedback and machine learning adaptation.

### v2

@@ 5. Evaluation Protocols and Datasets
In this section, we introduce the evaluation protocols and datasets for continul LLMs. In Section 5.1, we discuss common continual learning evaluation metrics adapted for this context, along with metrics designed specifically for continual LLMs. Then, in Section 5.2, we outline the datasets available for each discussed topic.
5.1. Evaluation Protocols
5.1.1. Interpreting Basic Continual Learning Evaluation Metrics in the Continual LLMs
In some literature, $\widetilde{\operatorname{OP}}$ is referred to as “example accuracy” (chen2024parameterizing, ), “whole accuracy” (song2023conpet, ), or “edit success rate” in CMR (hartvigsen2023aging, ). The concepts of Forgetting and Backward Transfer underpin various evaluation metrics, such as knowledge retention (jin2022lifelong, ), performance on unchanged knowledge (jang2022temporalwiki, ), average increased perplexity (AP+(qin2022elle, ), and test and edit retention rate in CMR (hartvigsen2023aging, ). We extend the notation of forward transfer in the vertical direction to represent the performance improvement on downstream tasks resulting from domain-adaptive pre-training (see Table 2). Forward Transfer is alternatively referred to as temporal generalization (jin2022lifelong, ) or knowledge transfer (lazaridou2021mind, ) in some literature.
5.1.2. Evaluation Metrics of Continual LLMs
@@
LAnguage Model Analysis (LAMA). LAnguage Model Analysis (LAMA) is an evaluation framework designed to probe the world knowledge embedded in language models (petroni2019language, ). It converts each world fact into a cloze statement, which is then inputted into the language models to predict the correct answer. LAMA has been extended for continual pre-training, particularly for those under the temporal shifts (jang2022temporalwiki, ; jang2022towards, ). In CKL, three LAMA benchmarks are constructed for different dimensions: InvariantLAMA assesses knowledge retention on time-invariant facts, UpdatedLAMA focuses on knowledge update, and NewLAMA evaluates knowledge acquisition (jang2022towards, ).
@@
Forgotten / (Updated + Acquired) Ratio (FUAR). As the performance of a pre-trained LLM is decomposed into a fine-grained set in CKL (jang2022towards, ), OP becomes a too general metric and cannot accurately reflect the balance and trade-offs of the model’s behavior. To address this issue, CKL proposes a joint evaluation metric FUAR (Forgotten / (Updated + Acquired) Ratio) for continual pre-training. A FUAR value of 1 represents an equal trade-off between the knowledge forgetting and knowledge learning: for each piece of updated or acquired knowledge, one piece of time-invariant knowledge is forgotten on average. A FUAR less than 1 suggests high learning efficacy, where more than one piece of knowledge is acquired at the expense of forgetting one piece of time-invariant knowledge.
@@
X-Delta. In TRACE (wang2023trace, ), the authors propose a set of “X-Delta” metrics for continual instruction tuning, quantifying the forward transfer on specific abilities of LLMs. Let’s denote a set of $M$ datasets $\{X_{1},X_{2},\cdots,X_{M}\}$ for task X. The baseline performances of the pre-trained LLM evaluated on these tasks are denoted as $\{b_{1}^{X},\cdots,b_{M}^{X}\}$. The model undergoes continuous fine-tuning on a different set of tasks, distinct from those used for evaluation. Throughout the sequential training process, the performance of the model after learning task $t$ on evaluation tasks $X_{i}$ is $R_{t,i}^{X}$. The X-Delta $\Delta R_{t}^{X}$ after learning task $t$ is defined as:
(14) $\Delta R_{t}^{X}$ $\triangleq\tfrac{1}{M}\sum_{m=1}^{M}(R_{t,i}^{X}-b_{i}^{X}).$
@@
In the public TRACE benchmark, the authors construct three sets of evaluation tasks to benchmark the ability of LLMs, including general ability, instruction following, and safety (wang2023trace, ).
@@
NLG Score. In continual model alignment, three prominent metrics used to evaluate different aspects of Natural language generation (NLG) are BLEU-4 (papineni2002bleu, ), METEOR (banerjee2005meteor, ), and ROUGE-L (lin2004rouge, ). BLEU-4(papineni2002bleu, ), designed for machine translation (MT), evaluates the precision of n-grams between the machine-generated and reference texts, focusing especially on four-word sequences to gauge fluency and adequacy. METEOR (banerjee2005meteor, ) also targets MT but aims to improve correlation with human judgment by considering synonyms and stemming, thus providing a more nuanced assessment of translation quality. On the other hand, ROUGE-L (lin2004rouge, ) is commonly applied in summarization tasks, assessing the longest common subsequence between the generated summary and a set of reference summaries, effectively measuring the recall of essential content. Each metric has its strengths and is tailored to specific kinds of language processing tasks, reflecting different dimensions of text generation quality.
5.2. Datasets
@@
In this section, we provide a comprehensive review of the datasets available for benchmarking continual LLMs, as illustrated in Table 4. We intentionally exclude datasets used for domain-adaptive pre-training LLMs in vertical domains such as legal, medical, and financial, unless they are specifically designed for continual domain-adaptive pre-training. Furthermore, we omit datasets used in general continual fine-tuning, as they have already been extensively studied in existing works (biesialska2020continual, ; ke2023continual, ).
@@
Datasets for Continual Pre-Training (CPT) and Domain Adaptive Pre-Training (DAP). Current research lacks a widely recognized benchmark for evaluating continual pre-training LLMs under temporal shifts. TimeLMs utilizes a series of Twitter corpora collected until 2022, sequentially pre-training RoBERTa models quarterly (loureiro2022timelms, ). CC-RecentNews, adopted as unlabeled pre-training data for LMs in CKL (jang2022towards, ), consists of recent news and serves as a single-stage dataset. Additionally, CKL introduces InvariantLAMA, NewLAMA, and UpdatedLAMA to assess the principles of continual knowledge learning. TWiki, a dataset derived from the articles of Wikipedia between August and December 2021, is curated and cleaned in TemporalWiki (jang2022temporalwiki, ). This dataset facilitates the exploration of incremental learning by providing the Diffsets between neighboring snapshots. For works that study the content-level distributional shifts in CPT and DAP, researchers often resort to a similar set of publicly available datasets (lo2020s2orc, ; xu2019bert, ; ni2019justifying, ) to construct their own test beds for continual learning algorithms. The DAPT dataset, developed by (gururangan2020dont, ), comprises four domains: BioMed and Computer Science from S2ORC (lo2020s2orc, ), News from (zellers2019defending, ), and Reviews from (he2016ups, ). In DAPT’s original study, each domain undergoes individual domain adaptive pre-training stages to demonstrate the universality of DAP’s effectiveness. Subsequent works, such as ELLE (qin2022elle, ) and Recyclable Tuning (qin2023recyclable, ), follow suit by employing these domains for multi-stage CPT. DEMix (gururangan2022demix, ) presents another large-scale dataset, featuring eight semantic domains with over 73.8 billion tokens. Alongside the training set, it includes eight additional datasets for validating the generalization ability of LLMs. On a smaller scale, CPT (ke2022continual-train, ) and DAS (ke2022continual-pre, ) datasets consist of four and eight domains, respectively, with approximately 3.12 million examples and a size of 4.16GB each. These datasets are constructed similarly to the aforementioned ones.
@@
Table 4. Summary of the existing benchmarks publicly available for Continual Learning LLMs. In the column of Name, we use the superscript “” to denote the lack of the dataset name and the name shown is that of the original paper. In this table, we deliberately omit the datasets used for domain-adaptive pre-training the vertical LLMs, as their main focus of development is not on continual learning. We also omit the datasets used for general continual fine-tuning, as they are extensively discussed in other existing surveys (biesialska2020continual, ; ke2023continual, ).
@@
TimeLMs (loureiro2022timelms, ) CPT Temporal Social Media 8 #Examples: 123.86M Tweets (loureiro2022timelms, ) code
CC-RecentNews (jang2022towards, ) CPT Temporal News 1 #Tokens: $\sim$168M Web (jang2022towards, ) code
@@
TWiki (jang2022temporalwiki, ) CPT Temporal General Knowledge 5 #Tokens: 4.7B Wikipedia (jang2022temporalwiki, ) code
@@
DAPT (gururangan2020dont, ) CPT DAP Content Multi-Domain 4 Size: 160GB BioMed (lo2020s2orc, ), CS (lo2020s2orc, ), News (zellers2019defending, ), Reviews (he2016ups, ) (gururangan2020dont, ) (qin2023recyclable, ) (qin2022elle, ) code
@@
CPT (ke2022continual-train, ) CPT Content Multi-Domain 4 #Examples: 3.12M Yelp (xu2019bert, ), S2ORC (lo2020s2orc, ), AG-News (zhang2015character, ) (ke2022continual-train, ) code
@@
DEMix (gururangan2022demix, ) CPT Content Multi-Domain 8 #Tokens: 73.8B 1B (chelba2014billion, ), CS (lo2020s2orc, ), Legal (caselaw2018, ), Med (lo2020s2orc, ) WebText (gokaslan2019OpenWeb, ), RealNews (zellers2019defending, ), Reddit (baumgartner2020pushshift, ), Reviews (ni2019justifying, ) (gururangan2022demix, ) code
@@
DAS (ke2022continual-pre, ) CPT DAP Content Multi-Domain 6 Size: 4.16GB Yelp (xu2019bert, ), Reviews (ni2019justifying, ), Papers (lo2020s2orc, ), PubMed (ke2022continual-pre, ) code
@@
SuperNI (wang2022supernaturalinstructions, ) CIT Content Mutli-Domain 16 #Tasks: 1616 #Examples: $\sim$5M GitHub (zhang2023citb, ; wang2024inscl, ) code
@@
CITB (zhang2023citb, ) CIT Content Mutli-Domain 19 #Tasks: 38 SuperNI (wang2022supernaturalinstructions, ) (zhang2023citb, ) code
@@
CoIN (chen2024coin, ) CIT Content Multi-Domain 8 #Examples: $\sim$1.14M RefCOCO (kazemzadeh-etal-2014-referitgame, ),RefCOCO+ (mao2016generation, ),RefCOCOg (mao2016generation, ) ImageNet (imagenet_cvpr09, ), VQAv2 (goyal2017making, ), ScienceQA (lu2022learn, ) TextVQA  (singh2019vqa, ), GQA (hudson2019gqa, ), VizWiz  (gurari2018vizwiz, ), OCR-VQA (mishraICDAR19, ) (chen2024coin, ) code
@@
TRACE (wang2023trace, ) CIT Content Mutli-Domain 8 #Examples: 56,000 ScienceQA (lu2022learn, ), FOMC (shah2023trillion, ), MeetingBank (hu2023meetingbank, ) C-STANCE (zhao-etal-2023-c, ), 20Minuten (kew-etal-2023-20, ), CodeXGLUE (lu2021codexglue, ), NumGLUE(mishra2022numglue, ) (wang2023trace, ) code
@@
NATURAL-INSTRUCTION (mishra2021natural, ) CIT Content Mutli-Domain 6 #Examples: 193k CosmosQA (huang2019cosmos, ), DROP (dua2019drop, ), Essential-Terms (khashabi-etal-2017-learning, ) MCTACO (zhou2019goingvacationtakeslonger, ), MultiRC (khashabi-etal-2018-looking, ), QASC (khot2020qasc, ) Quoref(dasigi-etal-2019-quoref, ) , ROPES (lin2019reasoning, ) , Winogrande (sakaguchi2019winogrande, ) (mishra2021natural, ) code
@@
IMDB (maas2011learning, ) CMA Content Social Media 1 Size: 217.35 MB IMDB (zhang2023copf, ) code
@@
HH-RLHF (bai2022training, ) CMA Content General Knowledge 1 Size: 28.1 MB Human Feedback (zhang2023copf, ) code
@@
Reddit TL;DR (volske2017tl, ) CMA Content Social Media 2 Size: 19.6 GB Reddit (zhang2023copf, ; zhangcppo, ) code
@@
Common Sense QA (lin2024mitigating, ) Reading Comprehension (lin2024mitigating, ) Translation (lin2024mitigating, ) CMA Content Multi-Domain 6 #Examples: $\sim$ 41.16M ARC Easy and Challenge (clark2018think, ), Race (lai2017race, ), PIQA (bisk2020piqa, ) SQuAD (rajpurkar2018know, ), DROP (dua2019drop, ) WMT 2014 French to English (bojar2014findings, ) (lin2024mitigating, ) see sources
@@
FEVER (fever, ) CMR Content General Knowledge 1 #Examples: 420k Wikipedia (de2021editing, ; hase2021language, ) code
@@
VitaminC (vitaminC, ) CMR Content General Knowledge 1 #Examples: 450k Wikipedia (mitchell2022memory, ) code
zsRE (zsRE, ) CMR Content General Knowledge 1 #Examples: 120M Wikireading (Wikireading, ) (hase2021language, ; meng2022locating, ; meng2022mass, ; hase2023does, ; hartvigsen2023aging, ; das2024larimar, ) -
@@
T-rex (T-rex, ) CMR Content General Knowledge 1 #Examples: 11M Dbpedia abstracts (Dbpedia, ) (li2022large, ; dong2022calibrating, ) code
@@
NQ (nq, ) CMR Content General Knowledge 1 #Examples: 320k Google queries, Wikipedia (hartvigsen2023aging, ) code
@@
CounterFact (meng2022locating, ) CMR Content General Knowledge 1 #Examples: 22k zsRE (zsRE, ) (meng2022locating, ; yu2023melo, ; hu2024wilke, ; das2024larimar, ) code
@@
SCOTUS (scotus, ) CMR Temporal Law 1 #Examples: 9.2k Supreme Court Database (hartvigsen2023aging, ) code
@@
Datasets for Continual Instruction Tuning. Measuring the effectiveness of CIT is crucial, particularly because traditional evaluation metrics may not be suitable for LLMs: many of them are overly simplistic and fail to comprehensively assess the model’s ability to learn continually. New benchmarks and metrics are required to evaluate both the retention of old knowledge and the integration of new instructions. TRACE (wang2023trace, ) stands as a continual learning benchmark designed specifically for LLMs, encompassing diverse tasks such as multilingual capabilities, code generation, and mathematical reasoning. CITB (zhang2023citb, ) represents another benchmark for CIT, incorporating both learning and evaluation protocols. It in addition demonstrates that replay generally yields the best performance across all methods. CoIN (chen2024coin, ) extends the benchmark to MLLMs, incorporating a balanced and diverse set of instructions from vision-language datasets.
@@
Datasets for Continual Model Refinement. Most datasets for continual model refinement can be categorized into two types (mazzia2023survey, ): fact checking and question answering. For fact checking, models are asked to verify the truthfulness of certain claims, typically modeled as a classification task. Key datasets include FEVER (fever, ) (used by (de2021editing, ; hase2021language, )) and VitaminC (vitaminC, ) (used by (mitchell2022memory, )), both sourced from Wikipedia. For question answering, models are tasked with providing specific answers instead of choices. Zero-shot Relation Extraction (zsRE) (zsRE, ) is the most widely employed dataset for this purpose  (hase2021language, ; meng2022locating, ; meng2022mass, ; hase2023does, ; hartvigsen2023aging, ; das2024larimar, ), alongside Natural Questions (NQ)  (nq, ) and T-rex  (T-rex, ). (meng2022locating, ) adapted zsRE with additional counterfactuals to create the more challenging CounterFact dataset, used by  (yu2023melo, ; hu2024wilke, ; das2024larimar, ). Beyond these two categories, SCOTUS (scotus, ) is also utilized (hartvigsen2023aging, ) in the assessment of continual model refinement through a document classification task for U.S. Supreme Court cases into 11 topics.
@@
Datasets for Continual Model Alignment. In the domain of reinforcement learning with human feedback (RLHF), several datasets are commonly employed across different studies to evaluate the adaptation and effectiveness of models under varying scenarios and continuous learning conditions. The IMDB (maas2011learning, ) and HH-RLHF (bai2022training, ) dataset, as introduced in (zhang2023copf, ) within their study on continual learning through optimal policy fitting, leverages data gathered from interactive RL scenarios to model human preferences dynamically. Similarly, the Reddit TL;DR dataset (volske2017tl, ) used by (zhangcppo, ; zhang2023copf, ) is focused on text summarization, providing a robust platform for testing the longevity and adaptability of learning algorithms under evolving conditions. Lastly, Common Sense QA (clark2018think, ; lai2017race, ; bisk2020piqa, ), Reading Comprehension (rajpurkar2018know, ; dua2019drop, ), and Translation (bojar2014findings, ), which are utilized in (lin2024mitigating, ) are selected to assess the challenges of aligning RL agents with human expectations without incurring significant performance penalties. Each of these datasets is pivotal in advancing the understanding of continual learning and the interplay between human feedback and machine learning adaptation.